Question 49
ProbabilityIn a recent zonal championship games involving 10 teams, teams X and Y were given probabilities 2/5 and 1/3 respectively of winning the gold in the football event. What is the probability that either team will win the gold?
- A.2/15
- B.7/15
- C.11/15
- D.13/15
Show the answer and explanation
Answer: C (11/15)
Explanation
Only one team can win, so add the probabilities: 2/5 + 1/3 = 6/15 + 5/15 = 11/15.